OpenStudy (anonymous):

Help pleeasee(:

OpenStudy (anonymous):

Determine which of the following statements is true concerning the values described in column #1 and column #2. The value found in column #1 is greater than the value found in column #2. The value found in column #1 is less than the value found in column #2. The value found in column #1 is equivalent to the value found in column #2. The relationship between column #1 and column #2 cannot be determined by the information given.

OpenStudy (anonymous):

Column 1-The x-coordinate of the vertex of the equation y = 2x2 − 4x + 12 Column 2- The x-coordinate of the vertex of the equation y = 4x2 + 8x + 3

OpenStudy (anonymous):

The one for column 1 is suppose to be 2x^2-4x+12

OpenStudy (anonymous):

@jim_thompson5910 @Nnesha @nincompoop @sleepyjess

OpenStudy (anonymous):

The vertex is found when y'=0 . So for column 1 y'=4x-4=0 . Then x=(4/4)=1 . Similarly you can solve for column 2, and you will get your answer.

Nnesha (nnesha):

use this formula to find vertex formula \[\huge\rm \frac{ -b }{ 2a }\]

Nnesha (nnesha):

quadratic equation \[\huge\rm Ax^2 + Bx + C= 0\]

Nnesha (nnesha):

y = 2x2 − 4x + 12 what is a and b in this quadratic equation ?

Nnesha (nnesha):

refresh!! if there are diamonds(question marks)

OpenStudy (anonymous):

I have no idea :/

Nnesha (nnesha):

quadratic equation \[\huge\rm Ax^2 + Bx + C= 0\] number at front of x^2 is a and a number at front of x is b

Nnesha (nnesha):

y = 2x2 − 4x + 12 what is a and b in this quadratic equation ? same with this one

OpenStudy (anonymous):

Oh i see, Sorry I didnt see you post it the first time. A would be 2 and b would be -4

Nnesha (nnesha):

yes right so plug that into the formula \[\huge\rm \frac{ -b }{ 2a }\]

OpenStudy (anonymous):

would the negative four stay negative?

Nnesha (nnesha):

b is negative and there is another negative in the formula so - times -4 = ?

OpenStudy (anonymous):

positive

Nnesha (nnesha):

yes right so what is final answer

OpenStudy (anonymous):

\[\frac{ 4 }{ 4}\]

Nnesha (nnesha):

looks good

OpenStudy (anonymous):

so 1

Nnesha (nnesha):

yep right now do same thing with 2nd Column use same formula substitute b and a into it and solve

OpenStudy (anonymous):

\[\frac{ 8 }{ 8 }\]

OpenStudy (anonymous):

once again 1

OpenStudy (anonymous):

so it would be C?

Nnesha (nnesha):

nope that not suppose to be one

OpenStudy (anonymous):

Gah.

Nnesha (nnesha):

 remeber there is negative sign and b is positive this tiem